OpenStudy (anonymous):

Solve the Equation, 3 y - = --------- y 9y-60

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (gw2011):

You can cross multiply this equation and you get: y^2=3(9y-60) y^2=27y-180 y^2-27y+180=0 (y-12)(y-15)=0 y-12=0 y=12 y-15=0 y=15 Therefore, y=12 and y=15